# A Lightweight and Gradient-Stable Neural Layer

### v3

@@ 5 Experiments on Larger Datasets
We conduct an investigation into the impact of the Han model on a widely used MNIST dataset lecun1998gradient and explored its robustness. We also evaluate the regression performance on five classic datasets, namely Bikesharing, Cal Housing, Elevators, Parkinsons, and Skillcraft fanaee2014event ; pace1997sparse ; Dua:2019 ; little2007exploiting ; thompson2013video . In addition, we assess the performance of the proposed model on four widely used image classification datasets, namely CIFAR10, CIFAR100, STL10, and Downsampled-ImageNet krizhevsky2009learning ; coates2011analysis ; chrabaszcz2017downsampled , where the last one is a downsampled version of ImageNet for affordability reasons.
All training is performed using PyTorch paszke2019pytorch on a shared cluster. Multiple runs are performed for each test instance, starting with different random initializations of the model parameters. All reported values are the average of at least five runs.
@@ 5.1.1 Same performance, fewer parameters
Our study aims to evaluate the model’s performance on MNIST dataset lecun1998gradient . In this study, we consider one hidden layer FCNet model, and $k$ Han-layer models, where $k$ ranges from 2 to 20. Notably, the width of each layer is fixed to the input size ($784$) in two models. To train our models, we use 10,000 instances for training and the remaining instances for testing. The details of other hyperparameters used in this study are presented in Table 4. We anneal the initial learning rate to 0.2 during half of the training period.
@@
Figure 8: The train/test performance in 1 hidden layer FCNet and $k$-layer HanNets ($k$ from 1 to 20).
@@ 5.1.2 Robustness by 1-Lipschitz
Many studies have established the robustness of 1-Lipschitz models against adversarial examples, for instance as evidenced by tsuzuku2018lipschitz . We investigate the contrast in robustness between Han and FC models using identical settings as in the previous subsection. The baseline models are one hidden layer FCNet with a width of 784, and a 20 hidden Han-layer model with the same width.
Adversarial examples are generated using PGD and CW attacks as detailed in carlini2017towards ; madry2018towards and summarized in Table 5. We extend our evaluation to defensive FC models trained with adversarial training and TRADES madry2018towards ; zhang2019theoretically using $\ell_{\infty}$ FGSM attack samples with a maximum perturbation of $\rho=0.1$, in addition to vanilla FCNet. To improve the robustness of HanNet, we adjust the value of weight decay in the optimizer and select the value from $\{0.01,0.05,0.1,0.2\}$ that offers the best robustness while maintaining clean accuracy above 0.9.
@@ 5.2 Regression
We conduct experiments on these five real-world regression datasets in Table 7. We choose Adam kingma2015adam optimizer which seems to be the method of choice for several works in that area including tsang2018neural in Table 8.
@@
We present test RMSE loss values and R-squared values in Table 9, where all values are averaged over 5 trials. Clearly, HanNet outperforms its “peer” FCNet1 by a notable margin, especially when fewer training samples are used, and is also better than FCNet2 which uses 15 times more parameters. We mention that the best test performance of HanNet is statistically the same as (or better than) that reported in tsang2018neural with an FCNet & NIT model that uses more than $800K$ parameters. Furthermore, with fewer parameters, HanNet appears less influenced by overfitting, as seen in Figure 9. Figures for all datasets in the three models are provided in Appendix.
Figure 9: The train (solid line) and test (dotted line) RMSE on Bikesharing. Left: $\delta=0.2$; right: $\delta=0.8$. Red: for FCNet1; blue: for HanNet.
@@ 5.3 Use of Han-layers in Image Classification
In this section, we experiment on integrating Han-layers into existing models for image classification. First, we use Han-layers to replace some MLP-layers in MLP-Mixer modelsmixertolstikhin2021 , and observe the resulting model performances. Next, we extend our investigation to a lightweight model, MobileViT mehta2021mobilevit .
@@ 5.3.1 Using Han-layers in MLP-Mixer
In literature, the term Multi-Layer Perceptron (MLP) is often used exchangeably with FCNet. Recently, MLP-dominated models have seen a wave of revivals for image recognition tasks mixertolstikhin2021 ; liu2021pay . MLP-dominated models (without multi-head attention) are much more concise than Transformer-based models dosovitskiy2020image ; touvron2021training but can still maintain test performances on very large-scale datasets. The motivation of MLP-Mixer mixertolstikhin2021 is to use the purely fully connected layers to remove attention architectures. An MLP-Mixer block is the elementary unit in MLP-Mixer models that consists of several FC-layers and skip-connections to form the map from input $X$ to output $Y$ as is shown in Figure 10.
@@
where we remove skip connections and layer normalizations (since HanNets does not suffer from gradient problems). The resulting Han/MLP-Mixer models are shown in Figure 11, where we arrange some Han-Mixer blocks after MLP-Mixer blocks (which may be empty). The main reason for us to combine the two types of Mixer blocks is that, short of drastically increasing network width, Han-Mixers alone cannot always provide enough model parameters for large-scale datasets. In addition, we use the convolution stem recommended by xiao2021early instead of the one in mixertolstikhin2021 .
@@ 5.3.2 Settings
We summarize various configurations of each block on different data sets in Table 10. We run Mixers using the following number of layers {1, 2, 4, 8, 12, 16} and select the best one on each dataset. We use the convolution stem recommended by xiao2021early instead of the one in mixertolstikhin2021 . Our convolutional stem designs use five layers, including four $3\times 3$ convolutions and a single $1\times 1$ final layer. The output channels are [64, 128, 256, 512, 512] on CIFAR10, CIFAR100 and STL10, and [128, 256, 512, 1024, 1024] on Downsampled-ImageNet. Some data augmentation methods are used, such as random HorizontalFlip, random augmentationcubuk2020randaugment , and Cutoutdevries2017improved . For these datasets, we choose a modified Adam called AdamW loshchilov2018decoupled . Detailed settings are in Table 11 below.
@@ 5.3.3 Results
Empirical evidence from MLP-based models mixertolstikhin2021 suggests that these models can achieve the state-of-the-art results on very large-scale datasets. Since the datasets used in our tests are not large enough to play to the full strengths of Mixer models, our aim here is not to observe how close our results can approach the state-of-the-art, but how Han-layer structures can impact the performance of Mixer models.
Table 12: Error rates (%) on STL10, CIFAR10, CIFAR100 and Downsampled-ImageNet datasets. On Downsampled-ImageNet, among MLPMixer ($i$) for $i=4,8,12,16$, $i=8$ gives the best results.
@@
WideResNet chrabaszcz2017downsampled 37.1 M 41.0 18.9
Table 12 reports test results from various MLP-Mixer and Han/MLP-Mixer models. On the smaller dataset STL10, HanMixers alone can already produce the best result. On larger datasets CIFAR10 and CIFAR100, enhanced performance from HanMixers is achieved either by adding a relatively few parameters (e.g., MLP (2) or (4) vs. MLP (2) or (4) + Han (16)), or even with far fewer parameters (e.g., MLP (4) vs. MLP (1)+Han (16)). On Downsampled-ImageNet, our Han/MLP-Mixer combination model clearly outperforms pure MLP-Mixers and offers a competitive performance with WideResNet chrabaszcz2017downsampled  (41% Top1 error) while using only 40% parameters. In summary, the benefits of using Han-Mixers should be empirically evident according to these experiments.
@@ 5.4 Using Han-layers in MobileViT
In this section, we incorporate Han-layers into a lightweight vision transformer model, MobileViT mehta2021mobilevit , and compare model performance after MLP-layers are replaced by Han-layers. MobileViT combines convolutional and fully connected layers to achieve good performance while being lightweight. Details of this model, which are beyond the scope of this paper, can be found in mehta2021mobilevit .
In MobileViT mehta2021mobilevit , an MLP-block is of the form
@@
which uses a layer normalization and Swish activation elfwing2018sigmoid . The above MLP-block is replaced, one-on-one, by the Han-block
@@
We experiment on two variants of the MobileViT model: MobileViT-XXS and MobileViT-XS with different sizes (see mehta2021mobilevit ). The training settings are similar to those in the previous MLP-Mixer experiments, except that we resize each image to $3\times 256\times 256$ according to the settings of MobileViT. Detailed settings are given in Table 13.
Table 13: Settings on the MobileViT experiments. Dim is the dimension of data vectors.
@@
CIFAR10/ CIFAR100 $3\times 256\times 256$ 0.001 0.1 128 300
@@
Testing performances of using MLP-blocks and Han-blocks are presented in Table 14. Across the three datasets (STL10, CIFAR10 and CIFAR100) and two model variants, the use of Han-blocks reduces the model parameter counts by approximately a half, while yielding comparable or better performance. These empirical findings again strongly support the use of Han-layers under suitable conditions.
Table 14: Error rates (%) on STL10, CIFAR10, CIFAR100 datasets, and on MobileViT-XXS and MobileViT-XS models.
@@ 6 Conclusions
We propose a lightweight layer structure (called Han-layer) for neural networks that has guaranteed gradient stability and 1-Lipschitz continuity. We provide extensive empirical evidence demonstrating the efficacy of Han-layers in reducing model parameters in certain neural networks. That is, when used strategically as substitutes for fully connected layers, a relatively small number of Han-layers can maintain or improve generalization performance. Moreover, we observe impressive generalization capabilities of HanNets on some synthetic datasets. These proof-of-concept results suggest that Han-layer can become a useful building block in constructing powerful deep neural networks. In the meantime, further investigations, especially theoretical ones, are still needed to fully assess the strength and weakness of Han-layers.
@@ 7 Acknowledgments
The work is supported by Shenzhen Research Institute of Big Data, and Shenzhen Science and Technology Program (Grant Number GXWD20201231105722002-20200901175001001)
@@ Appendix A Other Figures on Regression Datasets
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" alt="[LOGO]">

### v4

@@ 5 Experiments on Larger Datasets
We conduct an investigation into the impact of the Han model on a widely used MNIST dataset [36] and explored its robustness. We also evaluate the regression performance on five classic datasets, namely Bikesharing, Cal Housing, Elevators, Parkinsons, and Skillcraft [37, 38, 39, 40, 41]. In addition, we assess the performance of the proposed model on four widely used image classification datasets, namely CIFAR10, CIFAR100, STL10, and Downsampled-ImageNet [42, 43, 44], where the last one is a downsampled version of ImageNet for affordability reasons.
All training is performed using PyTorch [45] on a shared cluster. Multiple runs are performed for each test instance, starting with different random initializations of the model parameters. All reported values are the average of at least five runs.
@@ 5.1.1 Same performance, fewer parameters
Our study aims to evaluate the model’s performance on MNIST dataset [36]. In this study, we consider one hidden layer FCNet model, and $k$ Han-layer models, where $k$ ranges from 2 to 20. Notably, the width of each layer is fixed to the input size ($784$) in two models. To train our models, we use 10,000 instances for training and the remaining instances for testing. The details of other hyperparameters used in this study are presented in Table 4. We anneal the initial learning rate to 0.2 during half of the training period.
@@
Figure 8: The training (red) and testing (blue) performances in one hidden-layer FCNet and $k$-layer HanNets ($k$ from 1 to 20).
@@ 5.1.2 Robustness by 1-Lipschitz
Many studies have established the robustness of 1-Lipschitz models against adversarial examples, for instance as evidenced by [46]. We investigate the contrast in robustness between Han and FC models using identical settings as in the previous subsection. The baseline models are one hidden layer FCNet with a width of 784, and a 20 hidden Han-layer model with the same width.
Adversarial examples are generated using PGD and CW attacks as detailed in [47, 48] and summarized in Table 5. We extend our evaluation to defensive FC models trained with adversarial training and TRADES [48, 49] using $\ell_{\infty}$ FGSM attack samples with a maximum perturbation of $\rho=0.1$, in addition to vanilla FCNet. To improve the robustness of HanNet, we adjust the value of weight decay in the optimizer and select the value from $\{0.01,0.05,0.1,0.2\}$ that offers the best robustness while maintaining clean accuracy above 0.9.
@@ 5.2 Regression
We conduct experiments on these five real-world regression datasets in Table 7. We choose Adam [50] optimizer which seems to be the method of choice for several works in that area including [51] in Table 8.
@@
We present test RMSE loss values and R-squared values in Table 9, where all values are averaged over 5 trials. Clearly, HanNet outperforms its “peer” FCNet1 by a notable margin, especially when fewer training samples are used, and is also better than FCNet2 which uses 15 times more parameters. We mention that the best test performance of HanNet is statistically the same as (or better than) that reported in [51] with an FCNet & NIT model that uses more than $800K$ parameters. Furthermore, with fewer parameters, HanNet appears less influenced by overfitting, as seen in Figure 9. Figures for all datasets in the three models are provided in Appendix.
Refer to captionRefer to caption
Figure 9: The training (solid line) and testing (dotted line) RMSE on Bikesharing. Left: $\delta=0.2$; right: $\delta=0.8$. Red: for FCNet1; blue: for HanNet.
@@ 5.3 Use of Han-layers in Image Classification
In this section, we experiment on integrating Han-layers into existing models for image classification. First, we use Han-layers to replace some MLP-layers in MLP-Mixer models[52], and observe the resulting model performances. Next, we extend our investigation to a lightweight model, MobileViT [53].
@@ 5.3.1 Using Han-layers in MLP-Mixer
In literature, the term Multi-Layer Perceptron (MLP) is often used exchangeably with FCNet. Recently, MLP-dominated models have seen a wave of revivals for image recognition tasks [52, 54]. MLP-dominated models (without multi-head attention) are much more concise than Transformer-based models [55, 56] but can still maintain test performances on very large-scale datasets. The motivation of MLP-Mixer [52] is to use the purely fully connected layers to remove attention architectures. An MLP-Mixer block is the elementary unit in MLP-Mixer models that consists of several FC-layers and skip-connections to form the map from input $X$ to output $Y$ as is shown in Figure 10.
@@
where we remove skip connections and layer normalizations (since HanNets does not suffer from gradient problems). The resulting Han/MLP-Mixer models are shown in Figure 11, where we arrange some Han-Mixer blocks after MLP-Mixer blocks (which may be empty). The main reason for us to combine the two types of Mixer blocks is that, short of drastically increasing network width, Han-Mixers alone cannot always provide enough model parameters for large-scale datasets. In addition, we use the convolution stem recommended by [57] instead of the one in [52].
@@ 5.3.2 Settings
We summarize various configurations of each block on different data sets in Table 10. We run Mixers using the following number of layers {1, 2, 4, 8, 12, 16} and select the best one on each dataset. We use the convolution stem recommended by [57] instead of the one in [52]. Our convolutional stem designs use five layers, including four $3\times 3$ convolutions and a single $1\times 1$ final layer. The output channels are [64, 128, 256, 512, 512] on CIFAR10, CIFAR100 and STL10, and [128, 256, 512, 1024, 1024] on Downsampled-ImageNet. Some data augmentation methods are used, such as random HorizontalFlip, random augmentation[58], and Cutout[59]. For these datasets, we choose a modified Adam called AdamW [60]. Detailed settings are in Table 11 below.
@@ 5.3.3 Results
Empirical evidence from MLP-based models [52] suggests that these models can achieve the state-of-the-art results on very large-scale datasets. Since the datasets used in our tests are not large enough to play to the full strengths of Mixer models, our aim here is not to observe how close our results can approach the state-of-the-art, but how Han-layer structures can impact the performance of Mixer models.
Table 12: Error rates (%) on four datasets: STL10, CIFAR10, CIFAR100 and Downsampled-ImageNet.
@@
WideResNet [44] 37.1 M 41.0 18.9
Table 12 reports test results from various MLP-Mixer and Han/MLP-Mixer models. On the smaller dataset STL10, HanMixers alone can already produce the best result. On larger datasets CIFAR10 and CIFAR100, enhanced performance from HanMixers is achieved either by adding a relatively few parameters (e.g., MLP (2) or (4) vs. MLP (2) or (4) + Han (16)), or even with far fewer parameters (e.g., MLP (4) vs. MLP (1)+Han (16)). On Downsampled-ImageNet, our Han/MLP-Mixer combination model clearly outperforms pure MLP-Mixers and offers a competitive performance with WideResNet [44] (41% Top1 error) while using only 40% parameters. In summary, the benefits of using Han-Mixers should be empirically evident according to these experiments.
@@ 5.4 Using Han-layers in MobileViT
In this section, we incorporate Han-layers into a lightweight vision transformer model, MobileViT [53], and compare model performance after MLP-layers are replaced by Han-layers. MobileViT combines convolutional and fully connected layers to achieve good performance while being lightweight. Details of this model, which are beyond the scope of this paper, can be found in [53].
In MobileViT [53], an MLP-block is of the form
@@
which uses a layer normalization and Swish activation [61]. The above MLP-block is replaced, one-on-one, by the Han-block
@@
We experiment on two variants of the MobileViT model: MobileViT-XXS and MobileViT-XS with different sizes (see [53]). The training settings are similar to those in the previous MLP-Mixer experiments, except that we resize each image to $3\times 256\times 256$ according to the settings of MobileViT. Detailed settings are given in Table 13.
Table 13: Experiment settings on MobileViT. Dim represents data dimension.
@@
CIFAR10/100 $3\times 256\times 256$ 0.001 0.1 128 300
@@
Testing performances of using MLP-blocks and Han-blocks are presented in Table 14. Across the three datasets (STL10, CIFAR10 and CIFAR100) and two model variants, the use of Han-blocks reduces the model parameter counts by approximately one half, while yielding comparable or better performance. These empirical findings again strongly support the use of Han-layers under suitable conditions.
Table 14: Error rates (%) on datasets STL10, CIFAR10 and CIFAR100, for models MobileViT-XXS and MobileViT-XS.
@@ 6 Conclusions
We propose a lightweight layer structure (called Han-layer) for neural networks that has guaranteed gradient stability and 1-Lipschitz continuity. We provide extensive empirical evidence demonstrating the efficacy of Han-layers in reducing model parameters in certain neural networks. That is, when used strategically as substitutes for fully connected layers, a relatively small number of Han-layers can maintain or improve generalization performance. Moreover, we observe impressive generalization capabilities of HanNets on some synthetic datasets. These proof-of-concept results suggest that Han-layer can potentially become a useful building block in constructing powerful deep neural networks. Of course, further investigations, especially theoretical ones, are still needed to fully assess the strength and weakness of Han-layers.
@@ 7 Acknowledgments
The work is supported by Shenzhen Research Institute of Big Data, Shenzhen Science and Technology Program (Grant Number GXWD20201231105722002-20200901175001001) and NSFC 12141108.